# autocorrelation function CFR channel frequency response

### v3

@@ IV-C Comparison with Other Works
The MPM parameter distributions that have been just presented are herein compared with those heuristically or empirically obtained in previous scientific works. For example the presented statistical analysis confirms the assumptions made in [21], which statistically extended the MPM in (1) relying on the initial idea presented in [10]. Hence, it was inferred in [21] that the path gains $g_{i}$ turn out to be log-normally distributed variables multiplied by a random flip sign. The main difference is found for the statistics of $N$, supposed to have a Poisson distribution in [21], but that results into a GEV one. This difference was already hypothesized in [21] and is also justified by the different considered databases. The path lengths, $d_{i}$, were assumed to be log-normally distributed in [21], while the mixture of a Weibull and GEV has been obtained in this work. Additionally, the characterization method proposed in this paper provides a statistics for the other parameters, i.e., $a_{0}$, $a_{1}$, and $A$.
Other results concerning the MPM parameters distribution are listed in [34](Tab. 1), that extends the model in (1) to the MIMO transmission. In this case, the path gains $g_{i}$ are assumed uniformly distributed in $[-1,1]$ and the attenuation coefficient $a_{1}$ turns out to be constant. While, $K$ is normally distributed and $a_{0}$ exhibits a shifted exponential distribution. Instead, the channel median $A$ is uniformly or exponentially distributed depending on the considered channels. The same results are also summarized in [35, Tab. 2.10].
@@ IV-D Relationship Between Parameters
The development of MPM-based models requires a complete characterization of the involved parameters. This includes both their marginal PDFs, which have given in the previous subsection, and their joint distribution, which are examined in this one.
First, the relation between the path length, $d_{i}$, and the number of dominant paths, $N$, is explored. To this end, the probability distribution of the path length conditioned on $N$, $P_{r}\left(d_{i}|N\right)$ is estimated. Obtained results are given in Fig. 6, which gives additional details on the already observed bimodal behavior of the EPDF of $d_{i}$. As seen, there are almost no paths with lengths in the range from $1000-3000$ m, approximately, except for the case where $N$ is very large, in which some paths with lengths in this range can be found. The values of $d_{i}$ are concentrated in the low length region and in the region around 3000 m, being (approximately) uniformly distributed from zero (or almost zero) in the low length region. Furthermore, the range of path lengths in the low region increases with the number of dominant paths.
@@
Next, the relation between the absolute value of the path gain and the path length is assessed by estimating the conditional PDF of $|g_{i}|$ given $d_{i}$, $f_{|g_{i}|}\left(|g_{i}|\big{|}d_{i}\right)$. Results are displayed in Fig. 7. Interestingly, $|g_{i}|$ tends to a uniform distribution in the interval $(0,1)$ for very low values of $d_{i}$. As the path length increases, the range of values of $|g_{i}|$ decreases until it tends, again, to a uniform one in the interval $(0,1)$ for path lengths around $3000$ m. There is a clearly noticeable gap for path lengths between $1000-2500$ (approximately), where the few existing paths have very low gains.
@@
Finally, the correlation between the path gains is explored. Since the decimation process described in Section III.III-A yields a different number of paths for each measured channel, the following procedure is employed to allow comparing results from different channels. In order to distinguish the MPM model parameters before the decimation process is applied from the ones resulting after it, let us denote the latter by using tilded symbols. Hence, the original set of path lengths and path gains are expressed as $d_{n}$ and $g_{n}$, respectively, with $10\leq n\leq N-$ 1, where $N=$2554, and their corresponding sets after the decimation process as $\tilde{d}_{i}$ and $\tilde{g}_{i}$, respectively, with $0\leq i\leq\tilde{N}$, where $\tilde{N}$ is the number of dominant paths after the decimation process. These magnitudes are related as
@@
V Conclusion
@@
The MPM is one of the most common top-down models adopted in the literature about PLC. This paper explores the parameterization of this model, which was originally proposed for low-voltage power distribution networks in the band up to 20 MHz, for the indoor broadband scenario in the band up to 80 MHz. Since the number of parameters that result from the extension of the MPM model to this context is extremely large, some of them have been imposed on the basis of common sense intuition, others, such as the attenuation parameters and the set of initial path lengths, have been chosen according to physical assumptions, and the remaining ones are analytically retrieved by means of a WLS fitting.
@@
Once the MPM parameters have been computed according to the proposed strategy, they are related to the characteristics of the actual CFRs like the frequency selectivity and the average channel gain. Suitable distributions that fit the empirical PDF of each parameter are also given. Finally, the relation between the most relevant parameters of the model is explored. Provided results will aid the development improved channels models for indoor broadband PLC.

### v4

@@ IV-C Comparison with Other Works
The MPM parameter distributions that have been just presented are herein compared with those heuristically or empirically obtained in previous scientific works. For example the presented statistical analysis confirms the assumptions made in [Tonello12], which statistically extended the MPM in (1) relying on the initial idea presented in [Tonello07bis]. Hence, it was inferred in [Tonello12] that the path gains $g_{i}$ turn out to be lognormally distributed variables multiplied by a random flip sign. The main difference is found for the statistics of $N$, supposed to have a Poisson distribution in [Tonello12], but that results into a GEV one. This difference was already hypothesized in [Tonello12] and is also justified by the different considered databases. The path lengths, $d_{i}$, were assumed to be lognormally distributed in [Tonello12], while the mixture of a Weibull and GEV has been obtained in this work. Additionally, the characterization method proposed in this paper provides a statistics for the other parameters, i.e., $a_{0}$, $a_{1}$, and $A$.
Other results concerning the MPM parameters distribution are listed in [Hashmat2011, Table 1], that extends the model in (1) to the MIMO transmission. In this case, the path gains $g_{i}$ are assumed uniformly distributed in [1,1][-1,1] and the attenuation coefficient $a_{1}$ turns out to be constant. While, $K$ is normally distributed and $a_{0}$ exhibits a shifted exponential distribution. Instead, the channel median $A$ is uniformly or exponentially distributed depending on the considered channels. The same results are also summarized in [Lampe16, Table 2.10].
@@ IV-D Relationship Between Parameters
The development of MPM-based models requires a complete characterization of the involved parameters. This includes both their marginal PDFs, which have given in the previous subsection, and their conditional distribution, which are examined in this one.
First, the relation between the path length, $d_{i}$, and the number of dominant paths, $N$, is explored. To this end, the probability distribution of the path length conditioned on $N$, $f\left(d_{i}|N\right)$ is estimated. Obtained results are given in Fig. 7, which gives additional details on the already observed bimodal behavior of the EPDF of $d_{i}$. As seen, there are almost no paths with lengths in the range from 100030001000-3000 m, approximately, except for the case where $N$ is very large, in which some paths with lengths in this range can be found. The values of $d_{i}$ are concentrated in the low length region and in the region around 3000 m, being (approximately) uniformly distributed from zero (or almost zero) in the low length region. Furthermore, the range of path lengths in the low region increases with the number of dominant paths.
@@
Next, the relation between the absolute value of the path gain and the path length is assessed by estimating the conditional PDF of $|g_{i}|$ given $d_{i}$, $f\left(|g_{i}|\big|d_{i}\right)$. Results are displayed in Fig. 8. Interestingly, $|g_{i}|$ tends to a uniform distribution in the interval $(0,1)$ for very low values of $d_{i}$. As the path length increases, the range of values of $|g_{i}|$ decreases until it tends, again, to a uniform one in the interval $(0,1)$ for path lengths around $3000$ m. There is a clearly noticeable gap for path lengths between 100025001000-2500 (approximately), where the few existing paths have very low gains.
@@
Finally, the correlation between the path gains is explored. Since the decimation process described in Section III-A yields a different number of paths for each measured channel, the following procedure is employed to allow comparing results from different channels. In order to distinguish the MPM model parameters before the decimation process is applied from the ones resulting after it, let us denote the latter by using tilded symbols. Hence, the original set of path lengths and path gains are expressed as $d_{n}$ and $g_{n}$, respectively, with $0\leq n\leq N-1$, where N=N=2554, and their corresponding sets after the decimation process as $\tilde{d}_{i}$ and $\tilde{g}_{i}$, respectively, with $0\leq i\leq\tilde{N}$, where $\tilde{N}$ is the number of dominant paths after the decimation process. These magnitudes are related as
@@
V Applicability of the study and influence on the performance of PLC systems
This section discusses the use of the statistical analysis of the MPM parameters given in this work to the development of a stochastic MPM-based channel generator and also explores the relation between the model parameters and the performance of indoor PLC systems.
V-A Application to the development of a stochastic MPM
While the development of a statistical MPM-based channel generator is beyond the scope of this work, information provided herein can be used as a starting point to generate realizations of CFRs according to the MPM. To this end, the steps given below can be followed:
  1. 1.
    Generate a realistic value of the average channel gain, $G\;\textrm{(dB)}$}. To this end, the EPDF of $G\;\textrm{(dB)}$} for the set of measured channels is given in Fig. 10 (a), along with the PDF of the distribution that gives the best fit, whose parameters are provided in Table VI.
  2. 2.
    Generate a realistic value of the delay spread. This can be obtained by leveraging the relation between the average channel gain and the delay spread depicted in Fig. 10 (b), where the logarithm is taken to avoid the heteroscedasticity in the regression line, which is given by
    $\ln(\widehat{\textrm{ds}}(\mu\textrm{s}))=\alpha+\beta\cdot G(\textrm{dB}),$ (17)
    with $\alpha=-1.7499$ and $\beta=-2.7630\cdot 10^{-2}$. The distribution of the residuals of the fitting in (17) is given in Table VI.
  3. 3.
    Generate the values of $a_{0}$, $N$ and $A$ using the expressions given in Table III.
  4. 4.
    Since $a_{1}$ appears to be uncorrelated with the remaining parameters of the model, it can be generated independently according to the distribution given in Table V.
Sections IV-B and IV-D have revealed useful information about the probability distributions of $d_{i}$ and $g_{i}$ and their relation to other parameters of the MPM. However, further research is required to generate realistic sequences of these magnitudes. Our preliminary analysis suggests that the distribution of $d_{i}$ conditioned on $N$ can be modelled as a mixture of the Pearson and GEV distributions. Nevertheless, a larger number of measured channels are required to obtain accurate estimates of the distribution parameters for high values of $N$. A mathematical fitting of the conditional PDF of $|g_{i}|$ given $d_{i}$, illustrated in Fig. 8, is also required. The generation of sequences of $d_{i}$ requires knowledge of their joint distributions, which seems to be unrealistically complex. However, at least their covariance is required to obtain a linear model.
Figure 10: (a) Empirical and fitted PDFs of the average channel gain; (b) scatter plot of the relation between the logarithm of the delay spread and the average channel gain.
Table VI: Fitting distributions of the average channel gain and of the residuals of the fitting in (17)
Channel
parameter
Distribution
Distribution
parameters
Average channel gain
$G$ (dB)
GEV
k=0.2984k=-0.2984
$\sigma=13.9104$
$\mu=-43.025$
Residuals of the fitting in (17)
Normal
$\mu=1.2854$
$\sigma=0.3328$
V-B Relation between the MPM parameters and the performance of indoor PLC systems
[…]
For the sake of simplicity, continuous OFDM symbol transmission is assumed. A noise model with Gaussian distribution and PSD given by $S_{N}(f)=a+bf^{c}$ (dBm/kHz), with $f$ in MHz, $a=-1.7542\cdot 10^{3}$}, $b=1.6594\cdot 10^{3}$} and $c=-4.1565\cdot 10^{-3}$} is used [Cortes23].
Fig. 11 depicts the scatter plot of the PHY bit rate versus three key parameters of the MPM: the effective number of paths, $N$; the attenuation coefficient, $a_{0}$; and the normalization coefficient, $A$. As expected, performance decreases as $N$ and $a_{0}$ increases, in the former case because of the increased frequency selectivity of the channel and the latter because of the increased attenuation. As shown in Table VII, in both cases the relation can be fitted by means of a third-order polynomial function, although the cubic and quadratic terms have lower influence in the case of $a_{0}$, where the relation is almost linear. Regarding the relation to the normalization coefficient displayed in Fig. 11 (c), the performance increases with $A$. This is coherent with the fact that $A=1/\underset{i}{\max}(|g_{i}|)$, and the greater the absolute value of the path gains, the larger the frequency selectivity of the channel response.
Figure 11: Scatter plot and fitting curve of the estimated PHY bit rate (Mbit/s) versus (a) the number of effective paths, $N$; (b) the attenuation coefficient $a_{0}$; (c) the normalization coefficient, $A$.
Table VII: Parameters of the fitting expressions displayed in Fig. 11.
Fitting expression $\bm{\alpha}$ $\bm{\beta}$ $\bm{\gamma}$ $\bm{\delta}$
$\widehat{R}\textrm{ (Mbit/s)}=\alpha+\beta\cdot N+\gamma\cdot N^{2}+\delta\cdot N^{3}$ $8.1687$ 2.7978-2.7978 $3.8227\cdot 10^{-3}$ $1.8133\cdot 10^{-6}$
$\widehat{R}\textrm{ (Mbit/s)}=\alpha+\beta\cdot a_{0}+\gamma\cdot a_{0}^{2}+\delta\cdot a_{0}^{3}$ $7.0758\cdot 10^{2}$ $1.5554\cdot 10^{5}$ $1.359\cdot 10^{8}$ $3.5222\cdot 10^{10}$
$\widehat{R}\textrm{ (Mbit/s)}=\alpha\cdot e^{\beta\cdot A}+\gamma$ $4.8969\cdot 10^{-2}$ 21.0197-21.0197 $6.9578\cdot 10^{2}$ -
VI Conclusion
@@
The MPM is one of the most common top-down models adopted in the literature about PLC. This paper explores the parameterization of this model, which in the PLC context was firstly applied to low-voltage power distribution networks in the band up to 20 MHz, for the indoor broadband scenario in the band up to 80 MHz. Since the number of parameters that result from the extension of the MPM model to this context is extremely large, some of them have been imposed on the basis of common sense intuition, others, such as the attenuation parameters and the set of initial path lengths, have been chosen according to physical assumptions, and the remaining ones are analytically retrieved by means of a WLS fitting.
@@
Once the MPM parameters have been computed according to the proposed strategy, they are related to the characteristics of the actual CFRs like the frequency selectivity and the average channel gain. Suitable distributions that fit the empirical PDF of each parameter are also given. The relation between the most relevant parameters of the model is explored.
Finally, the applicability of the provided results for the development of stochastic MPM models for indoor broadband PLC and the influence of the MPM parameters on the performance of PLC system are discussed.